The shortest side of a triangle measures 15 feet less then the longest side. if the third side is 6 feet shorter then the longest side and the perimeter is 102 feet, find the measures of the three sides of the triangle
Help please and thank you

Answer by fractalier(6550)

If we call the longest side, x, then the other sides are x-15 and x-6.
We add them together to get 102, or
x + x-15 + x-6 = 102
3x - 21 = 102
3x = 123
x = 41
x-15 = 26
x-6 = 35
So the sides are 41 feet, 35 feet, and 26 feet.